OpenStudy (anonymous):

Expand the logarithmic expression. log8 a/2

OpenStudy (anonymous):

is the base 8?

OpenStudy (anonymous):

log(a/b)=log(a)-log(b)

OpenStudy (anonymous):

yes the base is 8

OpenStudy (anonymous):

\[8^{x} = a/2\]\[x = \frac{ \log_{}\frac{ a }{ 2 } }{ \log_{}8 }\]

OpenStudy (anonymous):

@tcarroll010 read the assignment

OpenStudy (anonymous):

already did

OpenStudy (anonymous):

these are the answers i have A. log82 – log8a B. log8a – log82 C. –a log82 D.